OpenStudy (anonymous):

if one zero and polynomial 3x^-8x-2k-1is seven times the other.find the zeroes andvalue of k.

OpenStudy (anonymous):

Again, \(3x^-8x-2k-1 \) is not a polynomial.

OpenStudy (anonymous):

hw

OpenStudy (anonymous):

Is it \[3x^2 -8x-2k-1 \]

OpenStudy (anonymous):

If this is the case let r and 7 r be the roots. The sum of the roots is \[ - \frac b a =\frac 8 3 =8 r\\ r=\frac 1 3 \]

OpenStudy (anonymous):

The product of the roots is \[ \frac c a \\\frac{7}{3\ 3}=-\frac{1}{3} (2 k+1)\\ k= - \frac 5 3 \]